OpenStudy (anonymous):

simplify 18-4i rad 4

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

can you draw it out?

OpenStudy (anonymous):

\[18-4i \]

OpenStudy (anonymous):



OpenStudy (anonymous):

@jim_thompson5910 its just the last step of the problem and rad 4 is 2 but im just not sure if you just mutiple it to the 18

jimthompson5910 (jim_thompson5910):

it seems odd to me, but if it really is \(\Large 18-4i\sqrt{4}\) then \[\Large 18-4i\sqrt{4} = 18-4i*2 = 18-8i\]

jimthompson5910 (jim_thompson5910):

but I think something is missing here because that was a bit too easy for this level of math

OpenStudy (anonymous):

the whole problem is (10-rad -72) + (8+rad -8) but i just dont know what to do for the last step @jim_thompson5910

OpenStudy (anonymous):

also your response isn't showing up for my computer could you type it?

jimthompson5910 (jim_thompson5910):

how is it showing up? as question marks or something?

OpenStudy (anonymous):

it says [Math Processing Error]

jimthompson5910 (jim_thompson5910):

ok one sec

jimthompson5910 (jim_thompson5910):

try this out

OpenStudy (anonymous):

got it! i multipled the two radical 2's which was my problem. thanks!

jimthompson5910 (jim_thompson5910):

you're welcome